OpenStudy (anonymous):

18-16y=-312 78x-16y=408 A: elimination, (-); (12,-33) B: Substitution (12,33) C:Substitution (12,-33) D: Elimination (-) (12,33)

OpenStudy (superdavesuper):

hmm...is the 1st eqn 18-16y=-312 or 18x -16y=-312?

OpenStudy (anonymous):

18x

OpenStudy (superdavesuper):

:P hahahaa im anticipating UR typos lol so what happens when u add the two eqns together?

OpenStudy (anonymous):

96x-32y=96

OpenStudy (superdavesuper):

Oooo sorry my mistake - should be subtracting one from the other...sorry!

OpenStudy (anonymous):

-60x=-720

OpenStudy (anonymous):

then i divide

OpenStudy (anonymous):

which is 12

OpenStudy (superdavesuper):

:) see what i meant by u no needing any more help? lol

OpenStudy (anonymous):

im confused on what next

OpenStudy (superdavesuper):

cuz' u solve it by ELIMINATING y at first, it is by..... ;)

OpenStudy (anonymous):

33?

OpenStudy (anonymous):

18x - 16y = -312 78x - 16y = 408 --->(-1) -------------- 18x - 16y = -312 -78x + 16y = -408 (result of multiplying by -1) --------------add -60x = - 720 x = -720/-60 x = 12 now sub 12 in for x in either of the equations to find y 18x - 16y = -312 18(12) - 16y = -312 216 - 16y = -312 -16y = -312 - 216 -16y = -528 y = -528/-16 y = 33 elimination (12,33)

OpenStudy (superdavesuper):

next is put x=12 back into either one eqn to solve for y...nevermind u figured it out on ur own already :)

OpenStudy (superdavesuper):

thank u, kelliegirl - i was just trying not to give out the solution myself :) but ur proof is good!

OpenStudy (anonymous):

I should not have interrupted your explaining...you were doing good..sorry about that superdave